### v2

# Distance-coupling as an Approach to Position and Formation Control

Abstract
Studies of coordinated motion in autonomous vehicle groups have become a significant topic of interest in recent years. In this paper, we study the case when the agents are limited to distance measurements from predetermined reference points, or anchors. A novel approach, referred to as distance-coupling, is proposed for approximating agent positions from anchor distances. By taking the difference between squared distance functions, we cancel out the quadratic term and obtain a linear function of the position for the point of interest. We apply this method to the homing problem and prove Lyapunov stability with and without anchor placement error; defining bounds on the region of attraction when the anchors are linearly transformed from their desired positions. To further increase complexity, we show how the policy can be implemented on the formation control problem for a set of autonomous agents. In this context, we prove the existence of the set of equilibria for the distance-coupled formation controller and make preliminary claims on its convergence. The performance for both controllers are subsequently validated through simulation.
I Introduction
Autonomously operated vehicles are commonly designed around high-cost measurement devices used for real-time localization. However, there exists instances, especially in low-cost systems, where the measurements available to the agents are limited to basic geometric information. We study the case when the information available is restricted to its Euclidean distance from $p$ fixed points, or anchors. Exploiting a linear, first-order model, we first address the homing problem. In this case, the primary goal is to move an agent from an arbitrary initial position to some predetermined equilibrium position using the available measurement data. Despite being based on nonlinear measurements, our controller is a simple linear function, and its stability can be proven in a straight-forward manner.
We also consider the problem of driving $m=p$ agents into some predetermined formation using only inter-agent distances. In this case, we show that the controller defined for the homing problem can be implemented with minimal adjustments. Subsequently, the set of equilibrium positions are proven through transformation invariance, and validated through simulation.
I-A Literature Review
Studies of distance-based position functions have primarily been investigated through barycentric coordinates [1]. These methods represent the positions of anchors as vertices of a polygon and utilize its geometry to derive the agent’s position. While this method is robust in many applications, the point is usually limited to the bounds of a convex polyhedron and maintains ill-defined behaviors when this assumption is broken. To overcome this, researchers have implemented policy switching schemes to the formation control problem with success in 2-D [2] and 3-D [3] coordinate frames.
For distance-based formation control, the most popular approach uses gradient-descent over a function defined on the measurement graph [4, 5]. The graph can take many forms (directed/undirected, fully connected, etc.), but the connections are predetermined and static. Papers by Lin, Dimarogonas and Choi are also among the first to demonstrate the stability of assorted gradient-based formation control policies through standard analysis methods [6, 7, 8].
Another noteworthy example is that of [9] which is among the first written on the topic and discusses a catalog of analytical, methods which can be used to stabilize formations within small perturbations of their equilibrium. Lastly, the papers [10, 11] discuss the formation control of second-order systems, and [12] defines a policy in terms of the complex Laplacian. The review paper [13] lays a good foundation for many of the methods available. In addition to formation control in a distance-based context, it also discusses methods for the position and displacement-based measurement cases.
I-B Paper Contributions
In this paper, we present a new method that algebraically cancels the nonlinear term from available distance functions. By taking the difference between squared-distance measurements, we show that we can accurately construct a linear approximation of the agent’s position in the environment. We refer to this method as distance-coupling.
The distance-coupled approach is novel in comparison to the literature as it does not have strict requirements on agent positions similar to those found in barycentric coordinate methods. Where preexisting methods primarily depend on the convexity of the anchor environment, our approach makes no such claims and allows the user to select anchor positions which are non-convex. However, we discuss degenerate collinearity cases and the minimum number of anchors required (Section III-A).
The new approach also enjoys better convergence expectations then the gradient-based alternative. For the homing problem, we prove that the linear controller exhibits global exponential convergence in the ideal case and a sufficiently large region of attraction when the anchors are linearly offset from their desired positions. For the formation control problem, we formally characterize the set of all equilibria through the controller’s transformation invariance, and confirm large regions of attraction empirically.
While we demonstrate our results in an $\mathbb{R}^{2}$ environment for simplicity, the methods discussed in Section III are generalized to $\mathbb{R}^{n}$.
II System Definition
We start by defining the agent dynamics and ideal controller, along with the measurements available to the system.
II-A System Dynamics
The agents are modeled as a linear, time-invariant single integrator of the form
$\dot{x}=u,$ (1)
where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{n}$ are the state and the control input, respectively. As common in the literature, we assume that the environment is unbounded and without obstacles. The goal of the homing problem is to move an agent from some initial condition to an equilibrium position denoted by $x^{(\text{eq})}$.
For the case that there are $m$-agents operating in a single environment, we can write (1) in matrix form such that
$\dot{X}=U,$ (2)
where $X\in\mathbb{R}^{n\times m}$ and $U\in\mathbb{R}^{n\times m}$ are the state and control for the set of agents defined by
$X=\begin{bmatrix}x_{1}&\cdots&x_{m}\end{bmatrix}\text{, and }U=\begin{bmatrix}u_{1}&\cdots&u_{m}\end{bmatrix}.$ (3)
II-B Ideal Controller
Since the system is linear, we define a simple state feedback controller that brings the agent to the goal position $x^{(\text{eq})}\in\mathbb{R}^{n}$ as
$u(x)=C(x-x^{(\text{eq})}),$ (4)
where $C\in\mathbb{R}^{n\times n}$ describes the policy coefficients. By standard linear system theory, if $C$ has eigenvalues with negative real parts, then the system is globally exponentially stable and the state will converge to the desired equilibrium from any initial condition [14].
Unfortunately, we are limited to the measurements described in the next section and cannot define a direct controller from the state of the agent. That said, (4) will be referred to as the ideal control.
II-C Anchor and Anchor-distance Sets
Figure 1: Anchor set where $p=3$ so that $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and $\mathcal{D}(x,A)=\{d(x,a_{1}),d(x,a_{2}),d(x,a_{3})\}$.
We assume the presence of $p$-anchors which are represented by the set
$\mathcal{A}=\{a_{i}\in\mathbb{R}^{n}:\forall i\leq p\},$ (5)
where $a_{i}$ is the position of the $i$-th anchor. We assume that the agent measures the distance from its position to each of the anchors. The anchor-distance set is thus defined as
$\mathcal{D}(x,\mathcal{A})=\{d(x,a_{i})\in\mathbb{R}:a_{i}\in\mathcal{A}\},$ (6)
where
$d(x,a_{i})=\sqrt{(x-a_{i})^{\intercal}(x-a_{i})}.$ (7)
An example for when $p=3$ is shown in Figure 1. In the interest of conciseness, the individual distance functions from the point $x$ are referred to as $d_{i}(x)=d(x,a_{i})$, and the anchor-distance set is referred to as $\mathcal{D}(x)=\mathcal{D}(x,\mathcal{A})$.
III Distance-coupled Position Control
In this section we show how (4) can be indirectly implemented as a function of the measurement set $\mathcal{D}(x)$ via distance-coupled functions. We then prove Lyapunov stability, characterize the region of attraction when the anchors are linearly transformed and show the behavior of the closed-loop system in simulation.
III-A Distance-coupled Position and Control
We start by squaring and expanding each element of the anchor-distance set (6) which results in
$d_{i}^{2}(x)$ $=(x-a_{i})^{\intercal}(x-a_{i})$ (8)
$=x^{\intercal}x-2x^{\intercal}a_{i}+a_{i}^{\intercal}a_{i}.$
The key step in our approach is to note that by subtracting the squared-distance functions for two anchors, the quadratic terms $x^{\intercal}x$ cancel out such that
$d_{i}^{2}(x)-d_{j}^{2}(x)=$ $\ (x^{\intercal}x-2x^{\intercal}a_{i}+a_{i}^{\intercal}a_{i})$ (9)
$-(x^{\intercal}x-2x^{\intercal}a_{j}+a_{j}^{\intercal}a_{j})$
$=$ $-2x^{\intercal}(a_{i}-a_{j})+a_{i}^{\intercal}a_{i}-a_{j}^{\intercal}a_{j}.$
Moving the state-related terms to the left we obtain
$-2(a_{i}-a_{j})^{\intercal}x=d_{i}^{2}(x)-d_{j}^{2}(x)-a_{i}^{\intercal}a_{i}+a_{j}^{\intercal}a_{j}.$ (10)
We reorganize the terms in (10) in matrix form with the intent of obtaining an expression for the full state $x$ through a linear system of equations:
$A$ $=\text{vstack}(\{-2(a_{i}-a_{j})^{\intercal}:\forall(i,j)\in\mathcal{I}\}),$ (11)
$b$ $=\text{vstack}(\{a_{i}^{\intercal}a_{i}-a_{j}^{\intercal}a_{j}:\forall(i,j)\in\mathcal{I}\}),$
$h(x)$ $=\text{vstack}(\{d_{i}^{2}(x)-d_{j}^{2}(x):\forall(i,j)\in\mathcal{I}\}),$
where $\mathcal{I}$ is the set of $(i,j)$ indices which make up every combination of the anchors included in (5). The vstack()($\cdot$)( ⋅ ) function simply converts a set of values into a vertically-oriented vector/matrix as appropriate. Using $|\mathcal{I}|$ to denote the cardinality of $\mathcal{I}$, we then have that $A\in\mathbb{R}^{|\mathcal{I}|\times n}$, and the vectors $b,h(x)\in\mathbb{R}^{|\mathcal{I}|}$.
Combining these with (10), the distance-coupled position of the agent can be computed by solving the linear system:
$Ax=h(x)-b$ (12)
$\Rightarrow A^{\intercal}Ax=A^{\intercal}(h(x)-b)$
[…]
where $K=(A^{\intercal}A)^{-1}A^{\intercal}$. Since the matrix $A^{\intercal}A$ is by definition positive definite, it is invertible if $A$ has full column rank with a subset of $n+1$ noncollinear anchor positions. While we do not include the proof here for conciseness, it is trivial and follows from standard linear algebra theory.
We refer to (12) as the distance-coupled position and can substitute it into (4) to obtain the corresponding distance-coupled controller:
$u(x)=C(K(h(x)-b)-x^{(\text{eq})}),$ (13)
where $u(x)$ is solely dependent on the sets $\mathcal{A}$ and $\mathcal{D}(x)$.
III-B Global Position Stability
We use Lyapunov’s method to prove stability of the distance-coupled controller (13) for an agent with model (1). Excluded from this paper for conciseness, a thorough definition of the Lyapunov stability criterion and the appropriate proofs can be found in [14]. We use $\mathbf{I}\in\mathbb{R}^{n\times n}$ to denote the identity matrix.
Proposition 1:
The distance-coupled control policy (13) is globally exponentially stable for an agent with dynamics (1) if $C=\alpha\mathbf{I}$ given that $\alpha<0$.
Proof.
Define the Lyapunov candidate using the inner product between $x$ and $x^{(\text{eq})}$:
$V(x)=(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})}).$ (14)
By definition, $V(x)=0$ at $x=x^{(\text{eq})}$ and is positive definite for all $x\neq x^{(\text{eq})}$, thus satisfying the first component to Lyapunov stability. Taking its derivative and substituting (13) for the control gives
$\dot{V}(x)$ $=2(x-x^{(\text{eq})})^{\intercal}\dot{x}=2(x-x^{(\text{eq})})^{\intercal}u(x)$ (15)
$=2(x-x^{(\text{eq})})^{\intercal}C(K(h(x)-b)-x^{(\text{eq})})$
Finally, using (12):
$\dot{V}(x)=2(x-x^{(\text{eq})})^{\intercal}C(x-x^{(\text{eq})}).$ (16)
Implying that $\dot{V}=0$ when $x=x^{(\text{eq})}$ and $\dot{V}<0$ given that $C=\alpha\mathbf{I}\prec 0$ with $\alpha<0$. Thus making the controller globally exponentially stable via Lyapunov stability [14].
Note that the proof can be extended to the case where $C$ is any stable matrix. However, we use the restriction $C=\alpha\mathbf{I}$ for the sake of analysis in the next section.
III-C Stability Under Transformations of the Anchor Set
Figure 2: Response of the distance-coupled controller (13) with $x^{(\text{eq})}$ at the origin and for varying magnitudes of randomly generated measurement noise, $\omega_{\varepsilon}$. The agents are given arbitrary initial positions around the origin.
Let us now study the case when the set $\mathcal{A}$ is transformed by some orthogonal rotation, $R\in\mathbb{R}^{n\times n}$, and offset $r\in\mathbb{R}^{n}$, but the distance-coupled controller components (11) remain constant. That is, $\tilde{a}_{i}\in\mathbb{A}$ are the coordinates which were originally used to derive (11), where $\mathbb{A}$ is referred to as the anchor frame. The true positions are denoted $a_{i}\in\mathbb{W}$, where $\mathbb{W}$ is the world frame.
For a given anchor we can thus define the transformation
$a_{i}=R\tilde{a}_{i}+r:\tilde{a}_{i}\in\mathcal{A}.$ (17)
We differentiate between the two frames as the transformation of the set $\mathcal{A}$ creates error in the agent position defined by (12). That is,
$\tilde{x}$ $=K(h(x)-b)=R^{\intercal}(x-r),$ (18)
where $x\in\mathbb{W}$ represents the world frame position of the agent and $\tilde{x}\in\mathbb{A}$ is its approximation in the anchor frame. The distance-coupled control function similarly becomes
$\tilde{u}(x)$ $=C(K(h(x)-b)-x^{(\text{eq})})$ (19)
$=C(\tilde{x}-x^{(\text{eq})})=C(R^{\intercal}(x-r)-x^{(\text{eq})})$
$=R^{\intercal}C(x-(Rx^{(\text{eq})}+r))$
Likewise to Proposition 1, we can prove stability in terms of a range of acceptable values for the rotation and offset of the anchors. In order to make concrete claims on the stability of a system in $\mathbb{R}^{2}$, we identify the range of acceptable $\theta$ which make $R$ positive definite. This finding will be subsequently used in Proposition 2.
Lemma 1:
A 2-D rotation defined by $R(\theta)\in\mathbb{R}^{2\times 2}:\theta\in\mathbb{R}$ is positive definite for the range $\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})$ $\forall k\in\mathbb{Z}$ where $\mathbb{Z}$ is the set of integers.
Proof.
Define the matrix $R$ in terms of Rodrigues’ formula for 2-D rotations:
$R(\theta)=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}$ (20)
where $\theta$ is some scalar angle. For some arbitrary point, $x$, we can say $R\succ 0$ when $x^{\intercal}Rx>0$ for all $x\neq 0$. Expanding this condition we get that
$x^{\intercal}Rx$ $=[x_{1}\ x_{2}]\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}>0$ (21)
$=(x_{1}^{2}+x_{2}^{2})\cos(\theta)>0,$
letting us conclude that $R(\theta)$ is positive definite so long as $\cos(\theta)>0$. This is true for the ranges
$\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})\ \forall k\in\mathbb{Z},$ (22)
where $\mathbb{Z}$ is the set of integers.
By this, the stability candidate developed in Proposition 1 can be reformulated for the rotated anchor case and used to identify the region of attraction for (19).
Proposition 2:
[…]
Proof.
Start by assuming that the new equilibrium, $\tilde{x}^{(\text{eq})}\in\mathbb{R}^{2}$, takes an equivalent transformation as that defined over the set $\mathcal{A}$, making the new Lyapunov candidate:
$V(x)$ $=(x-\tilde{x}^{(\text{eq})})^{\intercal}(x-\tilde{x}^{(\text{eq})})$ (23)
$=(x-(Rx^{(\text{eq})}+r))^{\intercal}(x-(Rx^{(\text{eq})}+r)).$
We thus have $V(x)=0:x=\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$ and $V(x)>0$ otherwise. Taking the derivative we get
$\dot{V}(x)$ $=2(x-(Rx^{(\text{eq})}+r))^{\intercal}\tilde{u}(x)$ (24)
$=2(x-(Rx^{(\text{eq})}+r))^{\intercal}C(K(h(x)-b)-x^{(\text{eq})}),$
where $\tilde{u}(x)$ is the controller defined in (19). Thus giving
$\dot{V}(x)$ $=2(x-(Rx^{(\text{eq})}+r))^{\intercal}C(R^{\intercal}(x-r)-x^{(\text{eq})})$ (25)
$=2(x-(Rx^{(\text{eq})}+r))^{\intercal}R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$
By inspection, we can conclude that $\dot{V}(x)=0:x=\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$. Moreover, $\dot{V}(x)$ is negative definite so long as $R^{\intercal}C\prec 0$. If we assume $C\prec 0$ by Proposition 1, then, $R^{\intercal}C\prec 0$ if $R$ is defined as in Lemma 1.
Remark:
While Proposition 2 is limited to $\mathbb{R}^{2}$ for simplicity, we identify that, given $C\prec 0$, the final claim holds for any system in $\mathbb{R}^{n}$ given that $R$ is a positive definite orthogonal rotation in $\mathbb{R}^{n}$.
III-D Distance-coupled Homing Results
Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately for $x^{(0)}$ chosen outside $Rx^{(\text{eq})}+r$.
Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately for $x^{(0)}$ chosen outside $Rx^{(\text{eq})}+r$.
Figure 4: Ratio of $100$ tests diverged for $100$ rotations, defined by $R:\theta\in[-2\pi,2\pi]$, of the anchor positions in $\mathcal{A}$.
We now validate the response of the distance-coupled controller (13) with varying degrees of noise in the measurement terms. The claims made on the regions of attraction defined in (22) will also be demonstrated. Note that the control coefficients are defined as $C=\alpha\mathbf{I}:\alpha=-5$ in all cases.
Let us define $\omega_{\varepsilon}$ as a uniformly distributed random variable with expected values in the range $[-\varepsilon,\varepsilon]$. The measurement noise is thus described as
$d_{i}(x,\varepsilon)=d_{i}(x)+\omega_{\varepsilon}.$
Figure 2 shows the response of the system with $p=3$ anchors and for varying degrees of noise, $\varepsilon$. The first case shows the ideal measurement scenario. That is, $\varepsilon=0$ with no error in the positions of the anchors. We place $x^{(\text{eq})}$ at the origin for simplicity. In this case, it is clear that there is no deviation between the true position and the anchor frame approximation. Likewise, the Lyapunov function converges ($V(x)\rightarrow 0$) as predicted from Proposition 1.
With an incorporation of a relatively large degree of measurement error ($\varepsilon=5$ and $\varepsilon=10$) we observe considerable differences between the approximation of the position from the distance-coupled function and the true position of the agent. Nonetheless, the agent is still capable of converging to a region around the equilibrium from any point.
Next, we validate the claims made in Section III-C for transformations of the anchors without updating (11). We first define the rotation $R:\theta=\tfrac{\pi}{4}$ and show that the agents converge with a policy defined by $R^{\intercal}C$. In the translation case, we sample various $r$ as uniform translations of the anchor set. Both sets of results are shown in Figure 4.
To demonstrate the region of attraction with respect to the rotation, $R$, of the anchors we can look at the behavior of the system for various angles $\theta$. Figure 4 shows the divergence rate for an agent with initial conditions
$x^{(0)}=x^{(\text{eq})}+\omega_{\varepsilon},$
where $\varepsilon=0.001$ is intentionally selected to be very small. We then show the results for $100$ uniformly spaced $\theta\in[-2\pi,2\pi]$ which were simulated until the candidate function converged ($V(x)\rightarrow 0$) or diverged ($V(x)\rightarrow\infty$). Each value in Figure 4 is the result of $100$ initial conditions. The system clearly stabilizes for all $\theta$ defined by the bounds in (22), and diverges otherwise. For a more explicit example, Figure 5 shows a series of agents initialized with $\omega_{\varepsilon=2.5}$ and $R:\theta=\tfrac{3\pi}{4}$ diverging from the origin as expected from (22).
IV Distance-coupled Formation Control
Figure 5: Distance-coupled controller response with the rotation, $R:\theta=\tfrac{3\pi}{4}$, of the anchor set (controller diverges).
Using the distance-coupled homing controller (13), we will define a new policy which drives $m$-agents to some predetermined formation. The only measurements each agent can make is the Euclidean distance it stands from each of the other agents in the environment. The following sections discuss the new system’s structure, defines the distance-coupled formation controller and proves certain properties analytically before demonstrating said claims in simulation.
IV-A Distance-coupled Formation Control
We utilize $m=p$ agents, with positions denoted by (3) and an intended formation described by the set $\mathcal{A}$. By defining our formation in terms of $\mathcal{A}$ we can exploit the notation from (12) to approximate the position of the agents in the anchor frame (discussed in Section IV-B).
In this scenario, each agent acts as an anchor for the others and vice versa. Thus, for an agent, $x_{k}$, we have the agent-distance set, $\mathcal{D}(x_{k},X)$, which is defined similarly to (6). We note that, in this form, the $k$-th element of the agent-distance set is always equal to 00.
We define the corresponding list of desired positions as
$X^{(\text{eq})}=\begin{bmatrix}x^{(\text{eq})}_{1}&\cdots&x^{(\text{eq})}_{m}\end{bmatrix}=\text{hstack}(\mathcal{A}),$ (26)
where hstack()($\cdot$)( ⋅ ) is the horizontal alternative to vstack()($\cdot$)( ⋅ ). In order to utilize (12) for all agents, we expand the measurement vector from (11) to the matrix:
$H(X)=\begin{bmatrix}h(x_{1})&\cdots&h(x_{m})\end{bmatrix},$ (27)
where $h(x_{k})$ is the distance-coupled agent measurements derived from $\mathcal{D}(x_{k},X)$, i.e.
$h(x_{k})=\text{vstack}(\{d^{2}(x_{k},x_{i})-d^{2}(x_{k},x_{j}):\forall(i,j)\in\mathcal{I}\}).$ (28)
We similarly arrange the squared anchor difference vector from (11) as
$B=b\mathbf{1}_{m}^{\intercal},$ (29)
where $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of all ones. This is valid as the vector $b$ is equivalent for every agent in $X$. Using the matrix notations defined above we can reformulate (12) and (13) as
$X=K(H(X)-B),$ (30)
$U(X)=C(K(H(X)-B)-X^{(\text{eq})}),$ (31)
where $K$ is similarly unchanged from (12). Finally, we can note that the only parameter which differs between the agents is the intended equilibrium position, identified by $x^{(\text{eq})}_{k}$ in the matrix $X^{(\text{eq})}$.
IV-B Formation Equivalence in the Anchor Frame
Before further inspection, we define formation equivalence (or, congruence) between the two frames of reference, $\mathbb{A}$ and $\mathbb{W}$, using selected tools from Procustes analysis [15]. Similar to the offsets defined in Section III-C, we have the agent positions in the world frame, $X\in\mathbb{W}$, with corresponding anchor frame coordinates $\tilde{X}\in\mathbb{A}$.
Definition 1:
Procrustes analysis defines the two sets, $X$ and $\tilde{X}$, as congruent if
Ψ s.t. $\exists\Psi\text{ s.t. }X-\langle x\rangle=\Psi(\tilde{X}-\langle\tilde{x}\rangle),$ (32)
given the orthogonal rotation, $\Psi(X,\tilde{X})=R(\theta)$, and the centroid of the formation, $\langle x\rangle=\tfrac{1}{m}\Sigma_{k=1}^{m}x_{k}$.
In practice, we can calculate an approximation of $\Psi$ using the Kabsch algorithm [16]. The resulting matrix is the most optimal rotation which superposes the set $\tilde{X}$ onto $X$.
Another method of denoting the transformation (32) is to combine the constant terms on the left and right sides into a single linear translation which superposes the anchor frame formation onto the world frame. This turns (32) into
$X-\langle x\rangle=\Psi(\tilde{X}-\langle\tilde{x}\rangle)\;\Rightarrow\;X=\Psi\tilde{X}+\psi,$ (33)
where $\psi=\langle x\rangle-\Psi\langle\tilde{x}\rangle$. The later sections will primarily use the shorthand definition of formation error (33). It can be further noted that the uniform scaling component to Procrustes analysis is omitted.
In order to measure policy convergence in Section IV-E, we define the formation error between $X$ and $X^{(\text{eq})}$ as the $L_{2}$-norm of the function (33) on individual agents, i.e.
$W(X)=\Sigma_{k=1}^{m}\|(\Psi x_{k}+\psi-x^{(\text{eq})}_{k})\|_{2}.$ (34)
IV-C Transformation Invariance
In order to discuss the equilibria of (31) we define transformation invariance and prove its application to the distance function. These claims will be used to similarly prove the invariance of (30) and (31) in the next section.
Definition 2:
A function $f(x,y):x,y\in\mathbb{R}^{n}$ is transformation invariant if $f(x,y)$ is constant under equivalent rotations and translations of both $x$ and $y$. That is,
$f(Rx+r,Ry+r)=f(x,y),$ (35)
given that $R\in\mathbb{R}^{n\times n}$ is an orthogonal rotation and $r\in\mathbb{R}^{n}$.
Lemma 2:
The distance function,
$d(x,y)=\sqrt{(x-y)^{\intercal}(x-y)},$ (36)
is transformation invariant by Definition 2.
Proof.
Define an arbitrary orthogonal rotation, $R$, and translation, $r$, of the original states $x$ and $y$ such that
$x\rightarrow Rx+r\text{, and }y\rightarrow Ry+r.$ (37)
Taking the distance between the transformed points:
$d(Rx+r,Ry+r)=\\ \sqrt{(Rx+r-(Ry+r))^{\intercal}(Rx+r-(Ry+r))}\\ \begin{aligned} &=\sqrt{(x-y)^{\intercal}R^{\intercal}R(x-y)}\\ &=\sqrt{(x-y)^{\intercal}(x-y)},\end{aligned}$
which is equivalent to (36). Thus concluding that the distance function follows by Definition 2.
IV-D Equilibria of the Distance-coupled Formation Controller
Figure 6: Distance-coupled formation control response for an arbitrary formation pattern and with varying levels of randomly generated offsets from the equilibrium positions: $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon}$.
Using Lemma 2 we establish the invariance of the distance-coupled functions in $H(X)$. Subsequently, we will use this claim to derive the equilibria of (31).
Proposition 3:
The distance-coupled position (30) is transformation invariant for any congruent set of positions $X=\Psi\tilde{X}+\psi$.
Proof.
Let there exists an orthogonal rotation $\Psi$ and translation $\psi$ so that $X$ and $\tilde{X}$ are congruent. Identifying that the dependent variables in (30) are solely contained in $H(X)$, we isolate a single row of the measurement terms for an arbitrary agent, $h(x_{k})$. That is,
$h_{i,j}(x_{k})=d^{2}(x_{k},x_{i})-d^{2}(x_{k},x_{j}),$ (38)
for given indices $(i,j)$. Applying the transformation $(\Psi,\psi)$ to the entire formation, we get
$h_{i,j}(\Psi\tilde{x}_{k}+\psi)=d^{2}(\Psi$ $\tilde{x}_{k}+\psi,\Psi\tilde{x}_{i}+\psi)$ (39)
$-d^{2}(\Psi\tilde{x}_{k}+\psi,\Psi\tilde{x}_{j}+\psi).$
Identifying the transformation invariance of the distance function from Lemma 2:
$h_{i,j}(\Psi\tilde{x}_{k}+\psi)$ $=d^{2}(\tilde{x}_{k},\tilde{x}_{i})-d^{2}(\tilde{x}_{k},\tilde{x}_{j})$ (40)
$=h_{i,j}(\tilde{x}_{k}).$
Therefore, for a transformation on the entire agent group, (40) holds for all $(i,j)\in\mathcal{I}$ and for any agent, $x_{k}$. This implies that every element in the measurement matrix, $H(X)$, is also transformation invariant. Combining this with the position given by (30), we get that
$H(X)=H(\Psi\tilde{X}+\psi)=H(\tilde{X})$ (41)
[…]
Proving that (30) is transformation invariant.
Proposition 4:
The controller (31) is transformation invariant for any congruent $X$ and $\tilde{X}$.
Proof.
Let us assume that Proposition 3 is true for the rotation $\Psi$ and translation $\psi$. Then we have that
$U(X)$ $=C(K(H(\Psi\tilde{X}+\psi)-B)-X^{(\text{eq})})$ (42)
$=C(K(H(\tilde{X})-B)-X^{(\text{eq})})$
$=U(\tilde{X}).$
Implying that (31) is transformation invariant for any congruent $X$ and $\tilde{X}$.
We can then use Proposition 4 to identify the equilibria of the distance-coupled formation controller.
Corollary 1:
[…]
Proof.
Assume that $\tilde{X}=X^{(\text{eq})}$ such that $U(\tilde{X})=0$. Then, for the world frame positions, $X=\Psi\tilde{X}+\psi$, and by Proposition 4 we have that
$U(X)=U(\Psi\tilde{X}+\psi)=U(\tilde{X})=U(X^{(\text{eq})})=0.$ (43)
Proving that the positions $X$ are at an equilibrium given they are congruent to $\tilde{X}$.
It should be noted that, while Corollary 1 defines the equilibria of (31), it makes no claims on its stability. However, because of the similarities between (13) and (31), it can be reasonably deduced that, for agent positions $X$ which are not congruent to $\tilde{X}$, we get similar bounds on the region of attraction as those shown in (22). This is demonstrated in the results for small perturbations of the agents from the equilibrium for various rotations.
IV-E Distance-coupled Formation Results
Figure 7: Response of formation controller with offset $\omega_{\varepsilon=10}$ and for the two agent sets with positions denoted by $X_{2}=X_{1}+r$.
Figure 7: Response of formation controller with offset $\omega_{\varepsilon=10}$ and for the two agent sets with positions denoted by $X_{2}=X_{1}+r$.
Figure 8: Region of attraction results with $p=4$ and $100$ repetitions at each $R:\theta\in[-2\pi,2\pi]$ for a randomly generated offset with $\omega_{\varepsilon=0.1}$.
While we do not present a proof for the stability of the distance-coupled formation controller, we can demonstrate empirically its behavior under various initial conditions. We first show the response of the policy when the initial conditions are randomly offset around $X^{(\text{eq})}$. This will also help to validate our claims on the equilibrium positions defined in Corollary 1. The region of attraction will then be plotted similarly to Figure 4, and the results will be compared.
To validate the response of (31) we initialize the agent group such that
$X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon},$
with $\varepsilon=10$ and $\varepsilon=25$. The simulation is then iterated and the behavior of the policy is shown. The performance of the formation error function $W(X)$ is plotted over time for each test, showing that it converges in these two cases.
Figure 8 uses two sets of agents to validate the convergence of the formation controller for similarly perturbed formations. The first set, $X_{1}$, is the reference set and has initial conditions which are generated similar to the Figure 6 example with $\varepsilon=10$. The second set is linearly offset from $X_{1}$ by a constant $r$. As expected, the two sets follow the same trajectory and end in equilibrium positions which differ by the initial translation, $r$.
Because of their close relationship, we tested if the controller would behave similarly to rotations about the origin as it does in the homing problem. We thus modified the initial condition formula for the agent set such that
$X^{(0)}=R(X^{(\text{eq})}+\omega_{\varepsilon}),$
with $\varepsilon=0.1$ and for the same list of $\theta$ described in Section III-D. The simulation was then iterated until the controller converged ($W(X)\rightarrow 0$) or diverged ($W(X)\rightarrow\infty$), and the proportion of tests which diverged for $100$ samples are shown in Figure 8. As expected, there is a strong correlation to the bounds defined in (22), with more complicated interplay in the formation policy case. Figure 9 shows an example of the controller diverging with $R:\theta=\pi$ and $\omega_{\varepsilon=0.1}$.
V Conclusion
Figure 9: Divergence of (31) with initial conditions $x^{(0)}=Rx^{(\text{eq})}$ given $R:\theta=\pi$ and $\omega_{\varepsilon=0.1}$.
In this paper we define the distance-coupling method as a suitable approach to the position and formation control problems. In the homing case we prove stability and show that it is held for linear transformations of the anchor positions, even when the policy components (11) are not updated. The distance-coupled formation controller is also studied, and the equilibrium conditions are established under the definitions of congruence and transformation invariance. Claims are made on the general performance of the system under arbitrary transformations, and the resulting convergence properties are demonstrated empirically.
Generally speaking, the approach described here is framed using a single-integrator model (1) for simplicity. That said, the distance-coupled position functions (12) and (30) can be applied to any localization problem in which the available measurements are a function of the distance from known reference points. In the context of systems with more complex dynamics, the distance-coupled policies defined here can be used to calculate suitable waypoints for the agents to follow by a more appropriate controller.

### v3

@@
-Distance-coupling as an Approach to Position and Formation Control
+references.bib
 Abstract
-Studies of coordinated motion in autonomous vehicle groups have become a significant topic of interest in recent years.
-In this paper, we study the case when the agents are limited to distance measurements from predetermined reference points, or anchors.
-A novel approach, referred to as distance-coupling, is proposed for approximating agent positions from anchor distances.
-By taking the difference between squared distance functions, we cancel out the quadratic term and obtain a linear function of the position for the point of interest.
-We apply this method to the homing problem and prove Lyapunov stability with and without anchor placement error; defining bounds on the region of attraction when the anchors are linearly transformed from their desired positions.
-To further increase complexity, we show how the policy can be implemented on the formation control problem for a set of autonomous agents.
-In this context, we prove the existence of the set of equilibria for the distance-coupled formation controller and make preliminary claims on its convergence.
-The performance for both controllers are subsequently validated through simulation.
+In this letter, we study the case of autonomous agents which are required to move to some new position based solely on the distance measured from predetermined reference points, or anchors.
+A novel approach, referred to as distance-coupling, is proposed for calculating the agent’s position exclusively from differences between squared distance measurements.
+The key insight in our approach is that, in doing so, we cancel out the measurement’s quadratic term and obtain a function of position which is linear.
+We apply this method to the homing problem and prove Lyapunov stability with and without anchor placement error; identifying bounds on the region of attraction when the anchors are linearly transformed from their desired positions.
+As an application of the method, we show how the policy can be implemented for distributed formation control on a set of autonomous agents, proving the existence of the set of equilibria.
 I Introduction
 Autonomously operated vehicles are commonly designed around high-cost measurement devices used for real-time localization.
-However, there exists instances, especially in low-cost systems, where the measurements available to the agents are limited to basic geometric information.
-We study the case when the information available is restricted to its Euclidean distance from $p$ fixed points, or anchors.
-Exploiting a linear, first-order model, we first address the homing problem.
-In this case, the primary goal is to move an agent from an arbitrary initial position to some predetermined equilibrium position using the available measurement data.
-Despite being based on nonlinear measurements, our controller is a simple linear function, and its stability can be proven in a straight-forward manner.
+However, there exists instances, especially in low-cost systems, when the measurements available to the agents are limited to basic geometric information.
+We study the case where the agent’s information is restricted to the Euclidean distance from $p$ fixed points, or anchors.
+To demonstrate, we control a single agent with linear, first-order dynamics moving from an arbitrary initial position to an equilibrium position given by target reference distances.
+Despite being based on nonlinear measurements, our controller is a simple linear function, and its stability can be proven in a straightforward manner.
 We also consider the problem of driving $m=p$ agents into some predetermined formation using only inter-agent distances.
-In this case, we show that the controller defined for the homing problem can be implemented with minimal adjustments.
-Subsequently, the set of equilibrium positions are proven through transformation invariance, and validated through simulation.
+We show that the controller defined for the homing problem can be implemented with minimal adjustments.
+The set of equilibrium positions are then proven and validated empirically.
 I-A Literature Review
-Studies of distance-based position functions have primarily been investigated through barycentric coordinates [1].
+Studies of distance-based position functions have primarily been investigated through barycentric coordinates [floater_generalized_2015].
 These methods represent the positions of anchors as vertices of a polygon and utilize its geometry to derive the agent’s position.
 While this method is robust in many applications, the point is usually limited to the bounds of a convex polyhedron and maintains ill-defined behaviors when this assumption is broken.
-To overcome this, researchers have implemented policy switching schemes to the formation control problem with success in 2-D [2] and 3-D [3] coordinate frames.
-For distance-based formation control, the most popular approach uses gradient-descent over a function defined on the measurement graph [4, 5].
-The graph can take many forms (directed/undirected, fully connected, etc.), but the connections are predetermined and static.
-Papers by Lin, Dimarogonas and Choi are also among the first to demonstrate the stability of assorted gradient-based formation control policies through standard analysis methods [6, 7, 8].
-Another noteworthy example is that of [9] which is among the first written on the topic and discusses a catalog of analytical, methods which can be used to stabilize formations within small perturbations of their equilibrium.
-Lastly, the papers [10, 11] discuss the formation control of second-order systems, and [12] defines a policy in terms of the complex Laplacian.
-The review paper [13] lays a good foundation for many of the methods available.
+To overcome this, researchers have implemented policy switching schemes to the formation control problem with success in 2-D coordinate frames [fathian_distributed_2019].
+For the formation control problem, the most popular approach implements gradient-descent with an objective function defined on the measurement graph [dimarogonas_connection_2008, dimarogonas_leaderfollower_2009].
+The graph can take many forms (directed/undirected, fully connected, etc.), but the communication connections are fixed.
+Papers by Dimarogonas, and Choi are also among the first to demonstrate the stability of assorted gradient-based formation control policies through standard analysis methods [dimarogonas_stability_2008, choi_distance-based_2021].
+The papers [ren_distributed_2007, chen_formation_2022] discuss the formation control of second-order systems, and [lin_distributed_2014] defines a policy in terms of the complex Laplacian.
+The review paper [oh_survey_2015] lays a good foundation for many of the methods available.
 In addition to formation control in a distance-based context, it also discusses methods for the position and displacement-based measurement cases.
+It discusses that, many instances the feedback laws require bearings or relative positions, so they are not purely distance-based.
 I-B Paper Contributions
-In this paper, we present a new method that algebraically cancels the nonlinear term from available distance functions.
-By taking the difference between squared-distance measurements, we show that we can accurately construct a linear approximation of the agent’s position in the environment.
-We refer to this method as distance-coupling.
-The distance-coupled approach is novel in comparison to the literature as it does not have strict requirements on agent positions similar to those found in barycentric coordinate methods.
-Where preexisting methods primarily depend on the convexity of the anchor environment, our approach makes no such claims and allows the user to select anchor positions which are non-convex.
-However, we discuss degenerate collinearity cases and the minimum number of anchors required (Section III-A).
-The new approach also enjoys better convergence expectations then the gradient-based alternative.
-For the homing problem, we prove that the linear controller exhibits global exponential convergence in the ideal case and a sufficiently large region of attraction when the anchors are linearly offset from their desired positions.
+In this letter, we present a new localization method which we refer to as distance-coupling.
+By taking the difference between squared-distance measurements, we show that we can accurately construct a linear approximation of the agent’s position in an environment given we know the positions of the reference points.
+Where preexisting methods depend on the convexity of the anchor environment and either bearings, or relative position information, our approach makes no such claims and allows the user to select anchor positions which are non-convex.
+However, we discuss degenerate collinearity cases and the minimum number of anchors required.
+For the homing problem, we prove that the linear controller exhibits global exponential convergence in the ideal case and a large region of attraction when the anchors deviate from their expected positions.
 For the formation control problem, we formally characterize the set of all equilibria through the controller’s transformation invariance, and confirm large regions of attraction empirically.
-While we demonstrate our results in an $\mathbb{R}^{2}$ environment for simplicity, the methods discussed in Section III are generalized to $\mathbb{R}^{n}$.
+Finally, we demonstrate our results in an $\mathbb{R}^{2}$ environment; with methods trivially generalized to any dimension.
 II System Definition
-We start by defining the agent dynamics and ideal controller, along with the measurements available to the system.
+We start by defining the agent dynamics, and the measurements available to the system at each point in time.
 II-A System Dynamics
-The agents are modeled as a linear, time-invariant single integrator of the form
-$\dot{x}=u,$ (1)
-where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{n}$ are the state and the control input, respectively.
-As common in the literature, we assume that the environment is unbounded and without obstacles.
-The goal of the homing problem is to move an agent from some initial condition to an equilibrium position denoted by $x^{(\text{eq})}$.
-For the case that there are $m$-agents operating in a single environment, we can write (1) in matrix form such that
-$\dot{X}=U,$ (2)
-where $X\in\mathbb{R}^{n\times m}$ and $U\in\mathbb{R}^{n\times m}$ are the state and control for the set of agents defined by
-$X=\begin{bmatrix}x_{1}&\cdots&x_{m}\end{bmatrix}\text{, and }U=\begin{bmatrix}u_{1}&\cdots&u_{m}\end{bmatrix}.$ (3)
-II-B Ideal Controller
-Since the system is linear, we define a simple state feedback controller that brings the agent to the goal position $x^{(\text{eq})}\in\mathbb{R}^{n}$ as
-$u(x)=C(x-x^{(\text{eq})}),$ (4)
-where $C\in\mathbb{R}^{n\times n}$ describes the policy coefficients.
-By standard linear system theory, if $C$ has eigenvalues with negative real parts, then the system is globally exponentially stable and the state will converge to the desired equilibrium from any initial condition [14].
-Unfortunately, we are limited to the measurements described in the next section and cannot define a direct controller from the state of the agent.
-That said, (4) will be referred to as the ideal control.
-II-C Anchor and Anchor-distance Sets
-Figure 1: Anchor set where $p=3$ so that $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and $\mathcal{D}(x,A)=\{d(x,a_{1}),d(x,a_{2}),d(x,a_{3})\}$.
+The agents are modeled as a linear, time-invariant single integrator of the form $\dot{x}=u,$ where $x\in\mathbb{R}^{n}$ and $u\in\mathbb{R}^{n}$ are the state and the control input, respectively.
+As is common in the literature, we assume that the environment is unbounded, and without obstacles.
+We will represent an agent’s intended equilibrium position as $x^{(\text{eq})}$.
+For the case where there are $m$ agents operating in a single environment, we can rewrite $\dot{x}$ in matrix form such that $\dot{X}=U$, where $X\in\mathbb{R}^{n\times m}$ and $U\in\mathbb{R}^{n\times m}$ are the state and control for the set of agents defined by $X=[x_{1}\cdots x_{m}]$, and $U=[u_{1}\cdots u_{m}]$.
+II-B Anchor and Anchor-distance Sets
+Figure 1: Anchor set with $p=3$ such that $\mathcal{A}=\{a_{1},a_{2},a_{3}\}$ and $\mathcal{D}(x,A)=\{d(x,a_{1}),d(x,a_{2}),d(x,a_{3})\}$.
 We assume the presence of $p$-anchors which are represented by the set
-$\mathcal{A}=\{a_{i}\in\mathbb{R}^{n}:\forall i\leq p\},$ (5)
-where $a_{i}$ is the position of the $i$-th anchor. We assume that the agent measures the distance from its position to each of the anchors. The anchor-distance set is thus defined as
+$\mathcal{A}=\{a_{i}\in\mathbb{R}^{n}\ \forall i\leq p\},$ (1)
+where $a_{i}$ is the position of the $i$-th anchor. We also assume that the agent measures the distance from its position to each of the anchors. The anchor-distance set is thus defined as
 $\mathcal{D}(x,\mathcal{A})=\{d(x,a_{i})\in\mathbb{R}:a_{i}\in\mathcal{A}\},$ (2)
-where
-$d(x,a_{i})=\sqrt{(x-a_{i})^{\intercal}(x-a_{i})}.$ (7)
+where $d(x,a_{i})=\sqrt{(x-a_{i})^{\intercal}(x-a_{i})}.$
 An example for when $p=3$ is shown in Figure 1.
-In the interest of conciseness, the individual distance functions from the point $x$ are referred to as $d_{i}(x)=d(x,a_{i})$, and the anchor-distance set is referred to as $\mathcal{D}(x)=\mathcal{D}(x,\mathcal{A})$.
+In future sections, we will also refer to the anchor distances as $d_{i}(x)=d(x,a_{i})$ and the set $\mathcal{D}(x)=\mathcal{D}(x,\mathcal{A})$.
+Again, we stress that the only measurements available to an agent is the set $\mathcal{D}(x)$, and their (possibly inexact) location with respect to the global environment.
 III Distance-coupled Position Control
-In this section we show how (4) can be indirectly implemented as a function of the measurement set $\mathcal{D}(x)$ via distance-coupled functions.
-We then prove Lyapunov stability, characterize the region of attraction when the anchors are linearly transformed and show the behavior of the closed-loop system in simulation.
-III-A Distance-coupled Position and Control
-We start by squaring and expanding each element of the anchor-distance set (6) which results in
+In this section we show how a linear controller can be indirectly implemented as a function of the (nonlinear) measurement set $\mathcal{D}(x)$ via distance-coupled functions.
+We then prove Lyapunov stability, characterize the region of attraction when the anchors are linearly transformed, and show the behavior of the closed-loop system in simulation.
+III-A Distance-coupled Position Functions
+We start by squaring and expanding each element of the anchor-distance set (2):
 $d_{i}^{2}(x)$ $=(x-a_{i})^{\intercal}(x-a_{i})$ (3)
 $=x^{\intercal}x-2x^{\intercal}a_{i}+a_{i}^{\intercal}a_{i}.$
-The key step in our approach is to note that by subtracting the squared-distance functions for two anchors, the quadratic terms $x^{\intercal}x$ cancel out such that
+The key step in our approach is to note that by subtracting the squared-distance functions for two anchors, the nonlinear term $x^{\intercal}x$ cancels out such that
 $d_{i}^{2}(x)-d_{j}^{2}(x)=$ $\ (x^{\intercal}x-2x^{\intercal}a_{i}+a_{i}^{\intercal}a_{i})$ (4)
@@
 $-2(a_{i}-a_{j})^{\intercal}x=d_{i}^{2}(x)-d_{j}^{2}(x)-a_{i}^{\intercal}a_{i}+a_{j}^{\intercal}a_{j}.$ (5)
+Note that $a_{i}$, and $a_{j}$ are expected to be known, as opposed to $d_{i}$ and $d_{j}$ which are measured.
 We reorganize the terms in (5) in matrix form with the intent of obtaining an expression for the full state $x$ through a linear system of equations:
@@
 where $\mathcal{I}$ is the set of $(i,j)$ indices which make up every combination of the anchors included in (1).
-The vstack()($\cdot$)( ⋅ ) function simply converts a set of values into a vertically-oriented vector/matrix as appropriate.
-Using $|\mathcal{I}|$ to denote the cardinality of $\mathcal{I}$, we then have that $A\in\mathbb{R}^{|\mathcal{I}|\times n}$, and the vectors $b,h(x)\in\mathbb{R}^{|\mathcal{I}|}$.
-Combining these with (10), the distance-coupled position of the agent can be computed by solving the linear system:
-$Ax=h(x)-b$ (12)
-$\Rightarrow A^{\intercal}Ax=A^{\intercal}(h(x)-b)$
-[…]
+The vstack()($\cdot$)( ⋅ ) function concatenates a set of values into a vertically-oriented vector/matrix as appropriate.
+Using $|\mathcal{I}|$ to denote the cardinality of $\mathcal{I}$ we have the matrix $A\in\mathbb{R}^{|\mathcal{I}|\times n}$, and the vectors $b,h(x)\in\mathbb{R}^{|\mathcal{I}|}$.
+Combining (5) and (6), the distance-coupled position of the agent can be computed by solving the linear system:
+$Ax=h(x)-b\Rightarrow A^{\intercal}Ax=A^{\intercal}(h(x)-b)$ (7)
+$\Rightarrow x=K(h(x)-b),$
 where $K=(A^{\intercal}A)^{-1}A^{\intercal}$.
-Since the matrix $A^{\intercal}A$ is by definition positive definite, it is invertible if $A$ has full column rank with a subset of $n+1$ noncollinear anchor positions.
-While we do not include the proof here for conciseness, it is trivial and follows from standard linear algebra theory.
-We refer to (12) as the distance-coupled position and can substitute it into (4) to obtain the corresponding distance-coupled controller:
+Since the matrix $A^{\intercal}A$ is always positive definite, it is invertible if $A$ has full column rank.
+This condition is satisfied if there exists a subset of at least $n+1$ noncollinear anchor positions.
+We refer to (7) as the distance-coupled position. The corresponding distance-coupled controller is defined as
 $u(x)=C(K(h(x)-b)-x^{(\text{eq})}),$ (8)
-where $u(x)$ is solely dependent on the sets $\mathcal{A}$ and $\mathcal{D}(x)$.
+where $C$ is a stable controller gain matrix, and $u(x)$ is solely dependent on the sets $\mathcal{A}$ (known) and $\mathcal{D}(x)$ (measured).
 III-B Global Position Stability
-We use Lyapunov’s method to prove stability of the distance-coupled controller (13) for an agent with model (1).
-Excluded from this paper for conciseness, a thorough definition of the Lyapunov stability criterion and the appropriate proofs can be found in [14].
-We use $\mathbf{I}\in\mathbb{R}^{n\times n}$ to denote the identity matrix.
-Proposition 1:
-The distance-coupled control policy (13) is globally exponentially stable for an agent with dynamics (1) if $C=\alpha\mathbf{I}$ given that $\alpha<0$.
-Proof.
-Define the Lyapunov candidate using the inner product between $x$ and $x^{(\text{eq})}$:
-$V(x)=(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})}).$ (14)
-By definition, $V(x)=0$ at $x=x^{(\text{eq})}$ and is positive definite for all $x\neq x^{(\text{eq})}$, thus satisfying the first component to Lyapunov stability.
-Taking its derivative and substituting (13) for the control gives
-$\dot{V}(x)$ $=2(x-x^{(\text{eq})})^{\intercal}\dot{x}=2(x-x^{(\text{eq})})^{\intercal}u(x)$ (15)
-$=2(x-x^{(\text{eq})})^{\intercal}C(K(h(x)-b)-x^{(\text{eq})})$
-Finally, using (12):
-$\dot{V}(x)=2(x-x^{(\text{eq})})^{\intercal}C(x-x^{(\text{eq})}).$ (16)
-Implying that $\dot{V}=0$ when $x=x^{(\text{eq})}$ and $\dot{V}<0$ given that $C=\alpha\mathbf{I}\prec 0$ with $\alpha<0$.
-Thus making the controller globally exponentially stable via Lyapunov stability [14].
-Note that the proof can be extended to the case where $C$ is any stable matrix. However, we use the restriction $C=\alpha\mathbf{I}$ for the sake of analysis in the next section.
-III-C Stability Under Transformations of the Anchor Set
 Figure 2: Response of the distance-coupled controller (8) with $x^{(\text{eq})}$ at the origin and for varying magnitudes of randomly generated measurement noise, $\omega_{\varepsilon}$.
 The agents are given arbitrary initial positions around the origin.
-Let us now study the case when the set $\mathcal{A}$ is transformed by some orthogonal rotation, $R\in\mathbb{R}^{n\times n}$, and offset $r\in\mathbb{R}^{n}$, but the distance-coupled controller components (11) remain constant.
-That is, $\tilde{a}_{i}\in\mathbb{A}$ are the coordinates which were originally used to derive (11), where $\mathbb{A}$ is referred to as the anchor frame.
-The true positions are denoted $a_{i}\in\mathbb{W}$, where $\mathbb{W}$ is the world frame.
-For a given anchor we can thus define the transformation
-$a_{i}=R\tilde{a}_{i}+r:\tilde{a}_{i}\in\mathcal{A}.$ (17)
-We differentiate between the two frames as the transformation of the set $\mathcal{A}$ creates error in the agent position defined by (12). That is,
-$\tilde{x}$ $=K(h(x)-b)=R^{\intercal}(x-r),$ (18)
-where $x\in\mathbb{W}$ represents the world frame position of the agent and $\tilde{x}\in\mathbb{A}$ is its approximation in the anchor frame. The distance-coupled control function similarly becomes
-$\tilde{u}(x)$ $=C(K(h(x)-b)-x^{(\text{eq})})$ (19)
-$=C(\tilde{x}-x^{(\text{eq})})=C(R^{\intercal}(x-r)-x^{(\text{eq})})$
-$=R^{\intercal}C(x-(Rx^{(\text{eq})}+r))$
-Likewise to Proposition 1, we can prove stability in terms of a range of acceptable values for the rotation and offset of the anchors.
-In order to make concrete claims on the stability of a system in $\mathbb{R}^{2}$, we identify the range of acceptable $\theta$ which make $R$ positive definite.
-This finding will be subsequently used in Proposition 2.
+We use Lyapunov’s direct method to prove stability of the distance-coupled controller (8) for an agent with dynamics defined by the single integrator model [khalil_nonlinear_2001].
+Proposition 1:
+The distance-coupled control policy (8) is globally exponentially stable for an agent with dynamics $\dot{x}=u$ if $C=\alpha\mathbf{I}$ with $\alpha<0$, and the identity $\mathbf{I}$.
+Proof.
+Define $V(x)=(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})})$ as the positive definite Lyapunov candidate.
+Then, using $\dot{x}=u$ and controller (8) with $C=\alpha\mathbf{I}$ it is trivial to show that $\dot{V}=2\alpha(x-x^{(\text{eq})})^{\intercal}(x-x^{(\text{eq})})$.
+Implying that $\dot{V}=0$ when $x=x^{(\text{eq})}$ and otherwise $\dot{V}<0$ given that $\alpha<0$.
+Therefore, for all $\alpha<0$ the controller is globally exponentially stable via Lyapunov stability [khalil_nonlinear_2001].
+Remark:
+Note that Proposition 1 can be extended to the case where $C$ is any stable matrix. However, we use the restriction $C=\alpha\mathbf{I}$ for the sake of analysis in the next section.
+III-C Stability Under Transformations of the Anchor Set
+We define the agent frame as $\mathbb{A}$, and the world frame $\mathbb{W}$.
+The controller (8) uses anchor positions $\tilde{a}_{i}\in\mathbb{A}$, but the measurements depend on the position $a_{i}\in\mathbb{W}$.
+The two sets are related by an unknown orthogonal rotation $R$ and offset $r$ such that $a_{i}=R\tilde{a}_{i}+r$.
+We identify that the transformation of the set $\mathcal{A}$ creates error in the agent position defined by (7), that is, $\tilde{x}=K(h(x)-b)=R^{\intercal}(x-r)$, where $x\in\mathbb{W}$ represents the world frame position of the agent and $\tilde{x}\in\mathbb{A}$ is its approximation in the anchor frame.
+Similarly, the controller (8) becomes
+$\tilde{u}(x)$ $=C(K(h(x)-b)-x^{(\text{eq})}),$ (9)
+$=C(\tilde{x}-x^{(\text{eq})})=C(R^{\intercal}(x-r)-x^{(\text{eq})}),$
+$=R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$
+In order to make stability claims similar to those discussed in Proposition 1, we will limit our system to $\mathbb{R}^{2}$, and identify the range of rotations which make $R\succ 0$.
 Lemma 1:
-A 2-D rotation defined by $R(\theta)\in\mathbb{R}^{2\times 2}:\theta\in\mathbb{R}$ is positive definite for the range $\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})$ $\forall k\in\mathbb{Z}$ where $\mathbb{Z}$ is the set of integers.
+An orthogonal rotation $R\in\mathbb{R}^{2\times 2}$ defined by $\theta\in\mathbb{R}$ is positive definite in the ranges $\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})$, where $k\in\mathbb{Z}$ and $\mathbb{Z}$ is the set of integers.
 Proof.
-Define the matrix $R$ in terms of Rodrigues’ formula for 2-D rotations:
-$R(\theta)=\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}$ (20)
-where $\theta$ is some scalar angle. For some arbitrary point, $x$, we can say $R\succ 0$ when $x^{\intercal}Rx>0$ for all $x\neq 0$. Expanding this condition we get that
-$x^{\intercal}Rx$ $=[x_{1}\ x_{2}]\begin{bmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}>0$ (21)
-$=(x_{1}^{2}+x_{2}^{2})\cos(\theta)>0,$
-letting us conclude that $R(\theta)$ is positive definite so long as $\cos(\theta)>0$. This is true for the ranges
-$\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})\ \forall k\in\mathbb{Z},$ (22)
-where $\mathbb{Z}$ is the set of integers.
-By this, the stability candidate developed in Proposition 1 can be reformulated for the rotated anchor case and used to identify the region of attraction for (19).
+Define the orthogonal rotation $R$ in terms of Rodrigues’ formula for 2-D rotations where the angle $\theta\in\mathbb{R}$.
+For some arbitrary point, $x$, we can say $R\succ 0$ when $x^{\intercal}Rx>0$ for all $x\neq 0$.
+After expanding, we get that $(x_{1}^{2}+x_{2}^{2})\cos(\theta)>0$; letting us conclude that $R$ is positive definite so long as $\cos(\theta)>0$.
+This is true for the ranges
+$\theta\in(2\pi k-\tfrac{\pi}{2},2\pi k+\tfrac{\pi}{2})\text{ for all }k\in\mathbb{Z},$ (10)
+where $\mathbb{Z}$ is the set of all integers.
+Using Lemma 1, the stability candidate developed in Proposition 1 can be reformulated for the rotated anchor case and used to identify the region of attraction for (9).
 Proposition 2:
-[…]
+For a 2-D system where the set $\mathcal{A}$ has been transformed by an orthogonal rotation, $R$, and translation, $r$, the transformed distance-coupled controller (9) describes an exponentially stable policy with equilibrium $x=Rx^{(\text{eq})}+r$ for any $r$ and any $R$ such that $R^{\intercal}C\prec 0$ (where $C$ is defined in Proposition 1 and $R$ in Lemma 1).
 Proof.
-Start by assuming that the new equilibrium, $\tilde{x}^{(\text{eq})}\in\mathbb{R}^{2}$, takes an equivalent transformation as that defined over the set $\mathcal{A}$, making the new Lyapunov candidate:
+Define the transformed equilibrium $\tilde{x}^{(\text{eq})}\in\mathbb{R}^{2}$ so that $\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$. Placing this into our original Lyapunov candidate we have,
 $V(x)$ $=(x-\tilde{x}^{(\text{eq})})^{\intercal}(x-\tilde{x}^{(\text{eq})})$ (11)
-$=(x-(Rx^{(\text{eq})}+r))^{\intercal}(x-(Rx^{(\text{eq})}+r)).$
-We thus have $V(x)=0:x=\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$ and $V(x)>0$ otherwise. Taking the derivative we get
-$\dot{V}(x)$ $=2(x-(Rx^{(\text{eq})}+r))^{\intercal}\tilde{u}(x)$ (24)
-$=2(x-(Rx^{(\text{eq})}+r))^{\intercal}C(K(h(x)-b)-x^{(\text{eq})}),$
-where $\tilde{u}(x)$ is the controller defined in (19). Thus giving
-$\dot{V}(x)$ $=2(x-(Rx^{(\text{eq})}+r))^{\intercal}C(R^{\intercal}(x-r)-x^{(\text{eq})})$ (25)
-$=2(x-(Rx^{(\text{eq})}+r))^{\intercal}R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$
-By inspection, we can conclude that $\dot{V}(x)=0:x=\tilde{x}^{(\text{eq})}=Rx^{(\text{eq})}+r$.
-Moreover, $\dot{V}(x)$ is negative definite so long as $R^{\intercal}C\prec 0$.
-If we assume $C\prec 0$ by Proposition 1, then, $R^{\intercal}C\prec 0$ if $R$ is defined as in Lemma 1.
-Remark:
-While Proposition 2 is limited to $\mathbb{R}^{2}$ for simplicity, we identify that, given $C\prec 0$, the final claim holds for any system in $\mathbb{R}^{n}$ given that $R$ is a positive definite orthogonal rotation in $\mathbb{R}^{n}$.
+$=(x-(Rx^{(\text{eq})}+r))^{\intercal}(x-(Rx^{(\text{eq})}+r)),$
+where $V=0$ when $x=Rx^{(\text{eq})}+r$ and $V>0$ otherwise. Taking the derivative of $V$ using (9) we get that
+$\dot{V}(x)=2(x-(Rx^{(\text{eq})}+r))^{\intercal}R^{\intercal}C(x-(Rx^{(\text{eq})}+r)).$ (12)
+By inspection, we can conclude that $\dot{V}(x)=0$ when $x=Rx^{(\text{eq})}+r$, and that $\dot{V}\prec 0$ so long as $R^{\intercal}C\prec 0$.
+If we assume $C\prec 0$ by Proposition 1, and $R$ is selected by the bounds defined in Lemma 1, then $R^{\intercal}C\prec 0$ if.
 III-D Distance-coupled Homing Results
-Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately for $x^{(0)}$ chosen outside $Rx^{(\text{eq})}+r$.
-Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately for $x^{(0)}$ chosen outside $Rx^{(\text{eq})}+r$.
-Figure 4: Ratio of $100$ tests diverged for $100$ rotations, defined by $R:\theta\in[-2\pi,2\pi]$, of the anchor positions in $\mathcal{A}$.
-We now validate the response of the distance-coupled controller (13) with varying degrees of noise in the measurement terms.
+We now validate the response of the distance-coupled controller (8) with varying degrees of noise in the measurement terms and different angles $\theta$ between $\mathbb{A}$ and $\mathbb{W}$.
 The claims made on the regions of attraction defined in (10) will also be demonstrated.
-Note that the control coefficients are defined as $C=\alpha\mathbf{I}:\alpha=-5$ in all cases.
-Let us define $\omega_{\varepsilon}$ as a uniformly distributed random variable with expected values in the range $[-\varepsilon,\varepsilon]$. The measurement noise is thus described as
-$d_{i}(x,\varepsilon)=d_{i}(x)+\omega_{\varepsilon}.$
-Figure 2 shows the response of the system with $p=3$ anchors and for varying degrees of noise, $\varepsilon$.
-The first case shows the ideal measurement scenario.
-That is, $\varepsilon=0$ with no error in the positions of the anchors.
-We place $x^{(\text{eq})}$ at the origin for simplicity.
-In this case, it is clear that there is no deviation between the true position and the anchor frame approximation.
+Note that the control coefficients are defined as $C=\alpha\mathbf{I}$ where $\alpha=-5$.
+We define $\omega_{\varepsilon}\in[-\varepsilon,\varepsilon]$ as a uniformly distributed random variable.
+The measurement noise is thus described as $d_{i}(x,\varepsilon)=d_{i}(x)+\omega_{\varepsilon}$.
+Figure 2 shows the response of the system with $p=3$ anchors and varying $\varepsilon$.
+In all tests we place $x^{(\text{eq})}$ at the origin.
+The first case shows the ideal measurement scenario, $\varepsilon=0$.
+It is clear that there is no deviation between the true position and the anchor frame approximation.
 Likewise, the Lyapunov function converges ($V(x)\rightarrow 0$) as predicted from Proposition 1.
-With an incorporation of a relatively large degree of measurement error ($\varepsilon=5$ and $\varepsilon=10$) we observe considerable differences between the approximation of the position from the distance-coupled function and the true position of the agent.
-Nonetheless, the agent is still capable of converging to a region around the equilibrium from any point.
-Next, we validate the claims made in Section III-C for transformations of the anchors without updating (11).
-We first define the rotation $R:\theta=\tfrac{\pi}{4}$ and show that the agents converge with a policy defined by $R^{\intercal}C$.
-In the translation case, we sample various $r$ as uniform translations of the anchor set.
-Both sets of results are shown in Figure 4.
-To demonstrate the region of attraction with respect to the rotation, $R$, of the anchors we can look at the behavior of the system for various angles $\theta$.
-Figure 4 shows the divergence rate for an agent with initial conditions
-$x^{(0)}=x^{(\text{eq})}+\omega_{\varepsilon},$
-where $\varepsilon=0.001$ is intentionally selected to be very small.
-We then show the results for $100$ uniformly spaced $\theta\in[-2\pi,2\pi]$ which were simulated until the candidate function converged ($V(x)\rightarrow 0$) or diverged ($V(x)\rightarrow\infty$).
-Each value in Figure 4 is the result of $100$ initial conditions.
-The system clearly stabilizes for all $\theta$ defined by the bounds in (22), and diverges otherwise.
-For a more explicit example, Figure 5 shows a series of agents initialized with $\omega_{\varepsilon=2.5}$ and $R:\theta=\tfrac{3\pi}{4}$ diverging from the origin as expected from (22).
+With a relatively large degree of measurement error ($\varepsilon=1$ and $\varepsilon=2.5$) we observe considerable differences between the anchor frame approximation and the true position.
+That said, the agent still converges to a region around the equilibrium from any point.
+Next, we validate the claims made in Section III-C for transformations of the anchors without updating the policy components (6).
+We first define the rotation $R$ with $\theta=\tfrac{\pi}{4}$, and show that the agents converge with a policy defined by $R^{\intercal}C$.
+We then demonstrate the effects of various translations $r$ over the anchor set and show that $x\rightarrow x^{(\text{eq})}+r$ as the simulation progresses.
+Both cases are shown in Figure 3.
+To demonstrate the instability defined in Proposition 2 we rotate the anchors by $\theta=\tfrac{3\pi}{4}$ which does not fall in the range defined in (10) with $k=0$ (Figure 4).
+Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately with (left) initial conditions around the origin, and (right) $x^{(0)}=x^{(\text{eq})}$.
+Figure 4: Distance-coupled controller response with the rotation, $R:\theta=\tfrac{3\pi}{4}$, of the anchor set (controller diverges).
 IV Distance-coupled Formation Control
-Figure 5: Distance-coupled controller response with the rotation, $R:\theta=\tfrac{3\pi}{4}$, of the anchor set (controller diverges).
 Using the distance-coupled homing controller (8), we will define a new policy which drives $m$-agents to some predetermined formation.
 The only measurements each agent can make is the Euclidean distance it stands from each of the other agents in the environment.
-The following sections discuss the new system’s structure, defines the distance-coupled formation controller and proves certain properties analytically before demonstrating said claims in simulation.
-IV-A Distance-coupled Formation Control
-We utilize $m=p$ agents, with positions denoted by (3) and an intended formation described by the set $\mathcal{A}$.
-By defining our formation in terms of $\mathcal{A}$ we can exploit the notation from (12) to approximate the position of the agents in the anchor frame (discussed in Section IV-B).
+The following sections discuss the new system’s structure, define the distance-coupled formation controller, and prove certain properties analytically before demonstrating in simulation.
+IV-A Distance-coupled Formation Functions
+We utilize $m=p$ agents, with positions denoted by $X$ and an intended formation described by the set $\mathcal{A}$.
+By defining our formation in terms of $\mathcal{A}$ we can use the notation from (7); approximating the positions of the agents in the anchor frame as discussed in Section IV-B.
 In this scenario, each agent acts as an anchor for the others and vice versa.
-Thus, for an agent, $x_{k}$, we have the agent-distance set, $\mathcal{D}(x_{k},X)$, which is defined similarly to (6).
-We note that, in this form, the $k$-th element of the agent-distance set is always equal to 00.
+Thus, for an agent $x_{k}$, we have the agent-distance set $\mathcal{D}(x_{k},X)$, defined similarly to (2).
+We note that the $k$-th element of $\mathcal{D}(x_{k},X)$ is always equal to 00 because of its inclusion in $X$.
 We define the corresponding list of desired positions as
 $X^{(\text{eq})}=\begin{bmatrix}x^{(\text{eq})}_{1}&\cdots&x^{(\text{eq})}_{m}\end{bmatrix}=\text{hstack}(\mathcal{A}),$ (13)
-where hstack()($\cdot$)( ⋅ ) is the horizontal alternative to vstack()($\cdot$)( ⋅ ). In order to utilize (12) for all agents, we expand the measurement vector from (11) to the matrix:
-$H(X)=\begin{bmatrix}h(x_{1})&\cdots&h(x_{m})\end{bmatrix},$ (27)
-where $h(x_{k})$ is the distance-coupled agent measurements derived from $\mathcal{D}(x_{k},X)$, i.e.
+where hstack()($\cdot$)( ⋅ ) is the horizontal concatenation function.
+In order to use (7) for all agents, we expand the measurement vector from (6) to the matrix $H(X)=[h(x_{1})\cdots h(x_{m})]$, where $h(x_{k})$ is the distance-coupled agent measurements derived from $\mathcal{D}(x_{k},X)$, i.e.
 $h(x_{k})=\text{vstack}(\{d^{2}(x_{k},x_{i})-d^{2}(x_{k},x_{j}):\forall(i,j)\in\mathcal{I}\}).$ (14)
-We similarly arrange the squared anchor difference vector from (11) as
-$B=b\mathbf{1}_{m}^{\intercal},$ (29)
-where $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of all ones.
+We similarly arrange the squared anchor difference vector from (6) as $B=b\mathbf{1}_{m}^{\intercal}$ where $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of all ones.
 This is valid as the vector $b$ is equivalent for every agent in $X$.
-Using the matrix notations defined above we can reformulate (12) and (13) as
+We can then reformulate (7) and (8) as
 $X=K(H(X)-B),$ (15)
 $U(X)=C(K(H(X)-B)-X^{(\text{eq})}),$ (16)
-where $K$ is similarly unchanged from (12).
-Finally, we can note that the only parameter which differs between the agents is the intended equilibrium position, identified by $x^{(\text{eq})}_{k}$ in the matrix $X^{(\text{eq})}$.
+where $K$ is unchanged from (7). Finally, we can note that the only parameter which differs between the agents is the intended equilibrium position.
 IV-B Formation Equivalence in the Anchor Frame
-Before further inspection, we define formation equivalence (or, congruence) between the two frames of reference, $\mathbb{A}$ and $\mathbb{W}$, using selected tools from Procustes analysis [15].
+Before further analysis, we define formation equivalence between the two frames of reference, $\mathbb{A}$ and $\mathbb{W}$, using tools from Procustes analysis.
 Similar to the offsets defined in Section III-C, we have the agent positions in the world frame, $X\in\mathbb{W}$, with corresponding anchor frame coordinates $\tilde{X}\in\mathbb{A}$.
 Definition 1:
-Procrustes analysis defines the two sets, $X$ and $\tilde{X}$, as congruent if
+Procrustes analysis defines the two sets, $X$ and $\tilde{X}$, as congruent if there exists a rotation $\Psi$ such that
 Ψ s.t. $\exists\Psi\text{ s.t. }X-\langle x\rangle=\Psi(\tilde{X}-\langle\tilde{x}\rangle),$ (17)
-given the orthogonal rotation, $\Psi(X,\tilde{X})=R(\theta)$, and the centroid of the formation, $\langle x\rangle=\tfrac{1}{m}\Sigma_{k=1}^{m}x_{k}$.
-In practice, we can calculate an approximation of $\Psi$ using the Kabsch algorithm [16]. The resulting matrix is the most optimal rotation which superposes the set $\tilde{X}$ onto $X$.
-Another method of denoting the transformation (32) is to combine the constant terms on the left and right sides into a single linear translation which superposes the anchor frame formation onto the world frame.
-This turns (32) into
-$X-\langle x\rangle=\Psi(\tilde{X}-\langle\tilde{x}\rangle)\;\Rightarrow\;X=\Psi\tilde{X}+\psi,$ (33)
-where $\psi=\langle x\rangle-\Psi\langle\tilde{x}\rangle$.
-The later sections will primarily use the shorthand definition of formation error (33).
-It can be further noted that the uniform scaling component to Procrustes analysis is omitted.
-In order to measure policy convergence in Section IV-E, we define the formation error between $X$ and $X^{(\text{eq})}$ as the $L_{2}$-norm of the function (33) on individual agents, i.e.
+where $\langle x\rangle=\tfrac{1}{m}\Sigma_{k=1}^{m}x_{k}$ is the centroid of the formation.
+In practice, we can compute an approximation of $\Psi$ using the Kabsch algorithm [kabsch_discussion_1978]. The resulting matrix is the optimal rotation to superpose the set $\tilde{X}$ onto $X$.
+The transformation (17) can also be written as $\Psi X+\psi$ where $\psi=\langle x\rangle-\Psi\langle\tilde{x}\rangle$.
+In order to measure convergence, we define the $L_{2}$-norm of the formation error as
 $W(X)=\Sigma_{k=1}^{m}\|(\Psi x_{k}+\psi-x^{(\text{eq})}_{k})\|_{2}.$ (18)
@@
 Definition 2:
-A function $f(x,y):x,y\in\mathbb{R}^{n}$ is transformation invariant if $f(x,y)$ is constant under equivalent rotations and translations of both $x$ and $y$. That is,
+A function $f(x,y):x,y\in\mathbb{R}^{n}$ is transformation invariant if $f(x,y)$ is constant under equivalent rotations and translations of both $x$ and $y$, that is,
 $f(Rx+r,Ry+r)=f(x,y),$ (19)
-given that $R\in\mathbb{R}^{n\times n}$ is an orthogonal rotation and $r\in\mathbb{R}^{n}$.
+for any orthogonal rotation $R\in\mathbb{R}^{n\times n}$ and offset $r\in\mathbb{R}^{n}$.
 Lemma 2:
@@
 Proof.
-Define an arbitrary orthogonal rotation, $R$, and translation, $r$, of the original states $x$ and $y$ such that
-$x\rightarrow Rx+r\text{, and }y\rightarrow Ry+r.$ (37)
-Taking the distance between the transformed points:
-$d(Rx+r,Ry+r)=\\ \sqrt{(Rx+r-(Ry+r))^{\intercal}(Rx+r-(Ry+r))}\\ \begin{aligned} &=\sqrt{(x-y)^{\intercal}R^{\intercal}R(x-y)}\\ &=\sqrt{(x-y)^{\intercal}(x-y)},\end{aligned}$
-which is equivalent to (36). Thus concluding that the distance function follows by Definition 2.
+Follows trivially from expanding and simplifying $d(Rx+r,Ry+r)$.
 IV-D Equilibria of the Distance-coupled Formation Controller
-Figure 6: Distance-coupled formation control response for an arbitrary formation pattern and with varying levels of randomly generated offsets from the equilibrium positions: $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon}$.
-Using Lemma 2 we establish the invariance of the distance-coupled functions in $H(X)$. Subsequently, we will use this claim to derive the equilibria of (31).
+Using Lemma 2 we establish the invariance of the distance-coupled functions in $H(X)$ before deriving the equilibria of (16).
 Proposition 3:
-The distance-coupled position (30) is transformation invariant for any congruent set of positions $X=\Psi\tilde{X}+\psi$.
+The distance-coupled position (15) is transformation invariant when $X=\Psi\tilde{X}+\psi$ are congruent.
 Proof.
-Let there exists an orthogonal rotation $\Psi$ and translation $\psi$ so that $X$ and $\tilde{X}$ are congruent.
-Identifying that the dependent variables in (30) are solely contained in $H(X)$, we isolate a single row of the measurement terms for an arbitrary agent, $h(x_{k})$.
-That is,
+Identifying that the dependent variables in (15) are solely contained in $H(x)$, we isolate a single row of the measurement terms for the $k$-th agent, $h(x_{k})$. This yields,
 $h_{i,j}(x_{k})=d^{2}(x_{k},x_{i})-d^{2}(x_{k},x_{j}),$ (21)
-for given indices $(i,j)$. Applying the transformation $(\Psi,\psi)$ to the entire formation, we get
-$h_{i,j}(\Psi\tilde{x}_{k}+\psi)=d^{2}(\Psi$ $\tilde{x}_{k}+\psi,\Psi\tilde{x}_{i}+\psi)$ (39)
-$-d^{2}(\Psi\tilde{x}_{k}+\psi,\Psi\tilde{x}_{j}+\psi).$
-Identifying the transformation invariance of the distance function from Lemma 2:
-$h_{i,j}(\Psi\tilde{x}_{k}+\psi)$ $=d^{2}(\tilde{x}_{k},\tilde{x}_{i})-d^{2}(\tilde{x}_{k},\tilde{x}_{j})$ (40)
-$=h_{i,j}(\tilde{x}_{k}).$
+for given indices $(i,j)$. Applying the transformation $x_{k}=\Psi\tilde{x}_{k}+\psi$ to each agent, and using Lemma 2, we get,
+$h_{i,j}(x_{k})$ $=h_{i,j}(\Psi\tilde{x}_{k}+\psi)$ (22)
+$=d^{2}(\tilde{x}_{k},\tilde{x}_{i})-d^{2}(\tilde{x}_{k},\tilde{x}_{j})=h_{i,j}(\tilde{x}_{k}).$
 Therefore, for a transformation on the entire agent group, (22) holds for all $(i,j)\in\mathcal{I}$ and for any agent, $x_{k}$.
@@
 […]
-Proving that (30) is transformation invariant.
+The claim then follows for (7).
 Proposition 4:
-The controller (31) is transformation invariant for any congruent $X$ and $\tilde{X}$.
+The controller (16) is transformation invariant when $X$ and $\tilde{X}$ are congruent.
 Proof.
-Let us assume that Proposition 3 is true for the rotation $\Psi$ and translation $\psi$. Then we have that
+By the same logic as Proposition 3, with rotation $\Psi$ and translation $\psi$, we have that
 $U(X)$ $=C(K(H(\Psi\tilde{X}+\psi)-B)-X^{(\text{eq})})$ (24)
-$=C(K(H(\tilde{X})-B)-X^{(\text{eq})})$
-$=U(\tilde{X}).$
+$=C(K(H(\tilde{X})-B)-X^{(\text{eq})})=U(\tilde{X}),$
 implying that (16) is transformation invariant for any congruent $X$ and $\tilde{X}$.
-We can then use Proposition 4 to identify the equilibria of the distance-coupled formation controller.
+We then use Proposition 4 to identify the equilibria of the distance-coupled formation controller.
 Corollary 1:
-[…]
+The formation controller (16) is at an equilibrium, $U(X)=0$, when $X$ is congruent to $X^{(\text{eq})}$.
 Proof.
-Assume that $\tilde{X}=X^{(\text{eq})}$ such that $U(\tilde{X})=0$. Then, for the world frame positions, $X=\Psi\tilde{X}+\psi$, and by Proposition 4 we have that
-$U(X)=U(\Psi\tilde{X}+\psi)=U(\tilde{X})=U(X^{(\text{eq})})=0.$ (43)
-Proving that the positions $X$ are at an equilibrium given they are congruent to $\tilde{X}$.
+Define $\tilde{X}$ at $X^{(\text{eq})}$ such that $U(\tilde{X})=0$.
+Then, by Proposition 4 we have that $U(X)=U(\Psi\tilde{X}+\psi)=U(\tilde{X})=0$.
+Thus, the formation $X=\Psi\tilde{X}+\psi$ is at an equilibrium given it is congruent to $X^{(\text{eq})}$.
 It should be noted that, while Corollary 1 defines the equilibria of (16), it makes no claims on its stability.
-However, because of the similarities between (13) and (31), it can be reasonably deduced that, for agent positions $X$ which are not congruent to $\tilde{X}$, we get similar bounds on the region of attraction as those shown in (22).
-This is demonstrated in the results for small perturbations of the agents from the equilibrium for various rotations.
+However, because of the similarities between (8) and (16), it can be reasonably concluded that, for agent positions $X$ which are not congruent to $\tilde{X}$, we get similar bounds on the region of attraction as those defined in Lemma 1.
+This claim is validated empirically in Section IV-E.
 IV-E Distance-coupled Formation Results
-Figure 7: Response of formation controller with offset $\omega_{\varepsilon=10}$ and for the two agent sets with positions denoted by $X_{2}=X_{1}+r$.
-Figure 7: Response of formation controller with offset $\omega_{\varepsilon=10}$ and for the two agent sets with positions denoted by $X_{2}=X_{1}+r$.
-Figure 8: Region of attraction results with $p=4$ and $100$ repetitions at each $R:\theta\in[-2\pi,2\pi]$ for a randomly generated offset with $\omega_{\varepsilon=0.1}$.
+Figure 5: Distance-coupled formation control response for an arbitrary formation and with varying levels of randomly generated offsets from the equilibrium positions: $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon=25}$.
+Figure 6: Response of formation controller with $X_{1}=X^{(\text{eq})}+\omega_{\varepsilon=10}$ and $X_{2}=X_{1}+r$.
+Figure 7: Region of attraction results with $p=4$ and $100$ repetitions at each $R$ with $\theta\in[-2\pi,2\pi]$ and for an initial position $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon=0.1}$.
 While we do not present a proof for the stability of the distance-coupled formation controller, we can demonstrate empirically its behavior under various initial conditions.
@@
 This will also help to validate our claims on the equilibrium positions defined in Corollary 1.
-The region of attraction will then be plotted similarly to Figure 4, and the results will be compared.
 To validate the response of (16) we initialize the agent group such that
-$X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon},$
-with $\varepsilon=10$ and $\varepsilon=25$.
+$X^{(0)}=R(X^{(\text{eq})}+\omega_{\varepsilon}),$
+with noise value $\varepsilon=25$, and rotation $R=\mathbf{I}$.
 The simulation is then iterated and the behavior of the policy is shown.
-The performance of the formation error function $W(X)$ is plotted over time for each test, showing that it converges in these two cases.
+The performance of the formation error function $W(X)$ is plotted over time; showing that it converges in this case.
 Figure 6 uses two sets of agents to validate the convergence of the formation controller for similarly perturbed formations.
 The first set, $X_{1}$, is the reference set and has initial conditions which are generated similar to the Figure 5 example with $\varepsilon=10$.
-The second set is linearly offset from $X_{1}$ by a constant $r$.
+The second set $X_{2}$ is linearly offset from $X_{1}$ by a constant $r$.
 As expected, the two sets follow the same trajectory and end in equilibrium positions which differ by the initial translation, $r$.
 Because of their close relationship, we tested if the controller would behave similarly to rotations about the origin as it does in the homing problem.
-We thus modified the initial condition formula for the agent set such that
-$X^{(0)}=R(X^{(\text{eq})}+\omega_{\varepsilon}),$
-with $\varepsilon=0.1$ and for the same list of $\theta$ described in Section III-D.
+We thus take $\varepsilon=0.1$ and select $R$ for a uniformly distributed set $\theta\in[-2\pi,2\pi]$.
 The simulation was then iterated until the controller converged ($W(X)\rightarrow 0$) or diverged ($W(X)\rightarrow\infty$), and the proportion of tests which diverged for $100$ samples are shown in Figure 7.
-As expected, there is a strong correlation to the bounds defined in (22), with more complicated interplay in the formation policy case.
-Figure 9 shows an example of the controller diverging with $R:\theta=\pi$ and $\omega_{\varepsilon=0.1}$.
+As conjectured, there is a strong correlation to the bounds defined in (10).
 V Conclusion
-Figure 9: Divergence of (31) with initial conditions $x^{(0)}=Rx^{(\text{eq})}$ given $R:\theta=\pi$ and $\omega_{\varepsilon=0.1}$.
 In this paper we define the distance-coupling method as a suitable approach to the position and formation control problems.
-In the homing case we prove stability and show that it is held for linear transformations of the anchor positions, even when the policy components (11) are not updated.
+In the homing case, we prove stability and show that it is held for linear transformations of the anchor positions, even when the policy components (6) are not updated.
 The distance-coupled formation controller is also studied, and the equilibrium conditions are established under the definitions of congruence and transformation invariance.
 Claims are made on the general performance of the system under arbitrary transformations, and the resulting convergence properties are demonstrated empirically.
-Generally speaking, the approach described here is framed using a single-integrator model (1) for simplicity.
+A key feature in our approach is that the controller can be implemented without relative position or bearings.
+Generally speaking, the approach described here is framed using a single integrator model for simplicity.
 That said, the distance-coupled position functions (7) and (15) can be applied to any localization problem in which the available measurements are a function of the distance from known reference points.
 In the context of systems with more complex dynamics, the distance-coupled policies defined here can be used to calculate suitable waypoints for the agents to follow by a more appropriate controller.
+[…]